OpenStudy (<3pandas18):

check my answer: 3x-3y=15 is equivilent to y=5-x?

OpenStudy (anonymous):

3x-3y=15 (1/3)(3x-3y+15) divide whole equation by 3 x-y=5 result x-x-y=5-x bring the x to the other side of the equation to leave y by itself -y=5-x result -(-y=5-x) multiply equation by -1 to have y a positive y=x-5 result

OpenStudy (anonymous):

y=5-x and y=x-5 are two different equations